OpenStudy (anonymous):

Simplify the rational expression. State any restrictions on the variable. n^4-11n^2+30/n^4-7n^2+10

OpenStudy (anonymous):

\[\frac{n^4-11n^2+30}{n^4-7n^2+10}=\frac{(n^2-5)(n^2-6)}{(n^2-5)(n^2-2)}=\frac{n^2-6}{n^2-2}\]

OpenStudy (anonymous):

I hope this helps...

OpenStudy (anonymous):

thank you, but which are the restrictions and how did you find them?

OpenStudy (anonymous):

well they are similar to a quadratic equation aren't they

OpenStudy (anonymous):

would it be n=/+-(squared)5, n=/(squared)2?

OpenStudy (anonymous):

oh ok u want to find the values of n, i thought u just said simplify

OpenStudy (anonymous):

it's both

OpenStudy (anonymous):

\[n= \pm \sqrt{6} \space and \space n= \pm 2\]

OpenStudy (anonymous):

it's not an option for my homework, it should have a 5 and 2

OpenStudy (anonymous):

it is impossible to have a 5 and 2 u'll have to have sqrt 5 and sqrt 2

OpenStudy (anonymous):

6

OpenStudy (anonymous):

i am pretty sure there is an error in the answers provided

OpenStudy (anonymous):

yeah I chose the answer that squared 5 and 2

OpenStudy (anonymous):

it should be 6 and 2, the 5s are canceled out

OpenStudy (anonymous):

as i showed u above

OpenStudy (anonymous):

there is no 5

OpenStudy (anonymous):

\[n= \pm \sqrt{6} \space and \space n= \pm \sqrt{2} \space or \space n^2= 6 \space or \space n^2= 2\]

OpenStudy (anonymous):

i think that i know what the problem is

OpenStudy (anonymous):

all of the answers have a =/ sign instead of an = sign

OpenStudy (anonymous):

\[\neq\]

OpenStudy (anonymous):

well is sqrt n a rational number or not?

OpenStudy (anonymous):

if it is not a rational number then u restrict it to not equal

OpenStudy (anonymous):

otherwise its equal

OpenStudy (anonymous):

i think this is clear to u now

OpenStudy (anonymous):

yes it does, thanks for clarifying =)

OpenStudy (anonymous):

all the best, so why is it TooMuch hehehe?

OpenStudy (anonymous):

because I have way too much to do all the time, anyway I'll give you best response

OpenStudy (anonymous):

thanks ur welcome...

OpenStudy (anonymous):

no problem, hopefully we can work together some time again :)

OpenStudy (anonymous):

my pleasure